# Fair Concurrent Training of Multiple Models in Federated Learning

### v1

@@ IV-A Fairness Guarantees
We first analyze the optimal solution to our fairness objective (2) to show that it reduces the variation in training accuracy, which intuitively improves fairness across tasks. Firstly, we denote $\textbf{w}^{*}_{s,\alpha}=G^{\alpha}=\min_{\textbf{w}_{s}}g^{\alpha}(\{\textbf{w}_{s}\})$, where
$G^{\alpha}=\min_{\mathbf{\textbf{w}_{s}}}\bigg{\{}\bigg{(}\frac{1}{S}\sum^{S}_{s=1}f^{\alpha}_{s}(\textbf{w}_{s})\bigg{)},\text{s.t.}\sum_{s}a_{k,s}=1\ \forall k\ \forall t\bigg{\}},$ (5)
where $a_{k,s}$ indicates that user $k$ trains task $s$ during epoch $t$ and $\sum_{s}a_{k,s}=1\ \forall k\ \forall t$ expresses the constraint that each client trains 1 task each epoch, and $\textbf{w}^{*}_{s,\alpha}$ is the global optimal solution when parameter $\alpha$ is used. In other words, $G^{\alpha}$ represents the optimal value of our MMFL objective (2).
To motivate the choice of $\alpha$, we show that for $\alpha=1,2$, a higher $\alpha$ puts more emphasis on fairness: we relate this $\alpha$-fairness metric to the well-known variance and cosine similarity metrics for similarity across task accuracies.
@@
$\alpha=2$ yields a smaller variance in the task accuracies than $\alpha=1$:
$\textbf{Var}(f_{1}(\textbf{w}_{1,2}^{*}),...,f_{n}(\textbf{w}_{n,2}^{*}))\leq\textbf{Var}(f_{1}(\textbf{w}_{1,1}^{*}),...,f_{n}(\textbf{w}_{n,1}^{*})).$
Proof Sketch: We followed the outline of the proof if [17]. As we are interested in task fairness, unlike [17], we replace client loss functions with the global task loss functions $f_{s}(\textbf{w}_{s,\alpha})$ in our derivations. We next combine the definitions of the variance and $\textbf{w}_{s,\alpha}^{*}$ in Eq. (5) to obtain our result. ∎
Next, we use the definition of cosine similarity to show that $\alpha=2$ results in a fairer solution than $\alpha=1$ (non-fairness).
@@
$\alpha=2$ results in a higher cosine similarity across tasks than $\alpha=1$, i.e.
$\frac{\frac{1}{S}\sum_{s=1}^{S}f_{s}(\textbf{w}_{s,2}^{*})}{\sqrt{\frac{1}{S}\sum_{s=1}^{S}f_{s}^{2}(\textbf{w}_{s,2}^{*})}}\geq\frac{\frac{1}{S}\sum_{s=1}^{S}f_{s}(\textbf{w}_{s,1}^{*})}{\sqrt{\frac{1}{S}\sum_{s=1}^{S}f_{s}^{2}(\textbf{w}_{s,1}^{*})}}.$ (6)
Proof Sketch: We followed the outline of [17]. Unlike [17], we are concerned with fairness across tasks, hence we replace user loss functions with the global task loss functions in our derivations. We further use the definition of cosine similarity, and the definition of $\textbf{w}_{s,\alpha}^{*}$ and $G^{\alpha}(\textbf{w}_{s})$ to obtain our result.∎
@@ IV-B Convergence Guarantees
We first characterize $B_{Sel}^{s}$, the probability that a specific set of clients $Sel$ is allocated (i.e., selected) for a specific task $s$ by our algorithm FedFairMMFL.
$B_{Sel}^{s}(\alpha)=\textstyle\left(\frac{f_{s}^{\alpha}(t)}{\sum_{s^{\prime}\in S}f_{s^{\prime}}^{\alpha}(t)}\right)^{|Sel|}\left(1-\frac{f_{s}^{\alpha}(t)}{\sum_{s^{\prime}\in S}f_{s^{\prime}}^{\alpha}(t)}\right)^{K-|Sel|}$ (7)
@@
$B_{Sel}^{s}(\alpha)$ depends on the losses of all tasks $s\in S$, and our $\alpha$-fair parameter $\alpha$. Next we introduce $\Gamma_{s}$, the gap between the loss evaluated at the global optimum $\mathbf{w}_{s}^{*}$, and the loss evaluated at client $k$’s optimal $\mathbf{w}^{*}_{k,s}$, for task $s$: $\Gamma_{s}=\sum_{k=1}^{K}p_{k}(F_{k,s}(\mathbf{w^{*}}_{s})-F_{k,s}(\mathbf{w}^{*}_{k,s}))=f_{s}^{*}-\sum_{k=1}^{K}p_{k}F_{k,s}^{*}.$ We further characterise the selection skew $\rho^{\alpha}(B^{s}_{Sel},\mathbf{w}^{\prime})$. The selection skew is a ratio of the expected ‘local-global objective gap’ of the selected clients under FedFairMMFL over the ‘local-global objective gap’ of all $K$ clients, where $\textbf{w}^{\prime}$ is the point at which $F_{k,s}$ and $f_{s}$ are evaluated. This skew depends on the expectation of client allocation combinations, and hence on $B_{Sel}^{s}$ and the $\alpha$-fair parameter $\alpha$:
$\rho^{\alpha}(B^{s}_{Sel},\mathbf{w}^{\prime})$
$=\frac{\sum_{Sel\in\mathbb{S}}{\color[rgb]{0,0,0}B_{Sel}^{s}(\alpha)}\sum_{k\in Sel}{\color[rgb]{0,0,0}\frac{p_{k}}{\sum_{k^{\prime}\in Sel}p_{k’}}}(F_{k,s}(\textbf{w}_{s}^{{}^{\prime}})-F_{k,s}^{*})}{f_{s}(\mathbf{w}^{\prime})-\sum_{k=1}^{K}p_{k}F_{k,s}^{*}}$
@@
We use $\overline{\rho}^{\alpha}$ and $\underline{\rho}^{\alpha}$ respectively to denote upper and lower bounds on the selection skew throuhgout the training. Recall that $\tau$ denotes the number of SGD epochs in local training. We then introduce the following assumptions used in our convergence analysis, which are standard in the FL literature [22].
Assumption 1: $F_{1,s},..,F_{k,s}$ are L-smooth, i.e. $F_{k,s}(\mathbf{v})\leq F_{k,s}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{k,s}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2},\ \forall\mathbf{v}$ and $\mathbf{w}$.
Assumption 2: $F_{1,s},..F_{k,s}$ are $\mu$-strongly convex, i.e. $F_{k,s}(\mathbf{v})\geq F_{k,s}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{k,s}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$, for all $\mathbf{v}$ and $\mathbf{w}$. Note that $\mu\leq L$ in from Assumption 1.
@@
Assumption 3: For all clients $k$, for the minibatch $\xi_{k,s}$ uniformly sampled at random, the stochastic gradient is unbiased: $\mathbb{E}[g_{k,s}(\mathbf{w}_{k,s},\xi_{k,s})]=\nabla F_{k,s}(\textbf{w}_{k,s})$, and its variance is bounded: $\mathbb{E}\|g_{k,s}(\mathbf{w}_{k,s},\xi_{k,s})]-\nabla F_{k,s}(\textbf{w}_{k,s})\|^{2}\leq\sigma^{2}$.
@@
Assumption 4: Bounded stochastic gradient. For all clients $k$, the expected square norm of the stochastic gradient is bounded $\mathbb{E}\|g_{k,s}(\textbf{w}_{k,s},\xi_{k,s})\|^{2}\leq G^{2}$.
@@
The expected average discrepancy between the global model for task $s$, $\overline{\textbf{w}}^{(t)}_{s}$ and client $k$’s model for $s$, $\textbf{w}_{k,s}^{(t)}$, under selection $k\in Sel(t)$, can be bounded as
$\mathbb{E}[\sum_{k\in Sel(t)}\frac{p_{k}}{\sum_{k^{\prime}\in Sel(t)}p_{k^{\prime}}}\|\overline{\textbf{w}}^{(t)}_{s}-\textbf{w}_{k,s}^{(t)}\|^{2}]$ (8)
$\leq 16\eta_{t}^{2}\tau^{2}G^{2}\sum_{Sel(t)\in\mathbb{S}(t)}B_{Sel(t)}^{s}(\alpha)P^{k,2}_{Sel(t)}$ (9)
where $P^{k,2}_{Sel(t)}=\sum_{\begin{subarray}{c}k\neq k^{\prime},\\ k,k^{\prime}\in Sel(t)\end{subarray}}p_{k}\frac{p_{k^{\prime}}}{(\sum_{k^{{}^{\prime\prime}}\in Sel(t)}p_{k^{\prime\prime}})^{2}}$.
In the next theorem, without loss of generality we show the convergence rate of a task in the MMFL setting under FedFairMMFL, and show its dependence on the relative loss levels of all tasks. When the prevailing loss level of task $s$ is higher, its convergence will be sped up under our algorithm.
@@
Suppose we use learning rate $\eta_{t}$ at each round $t$. Under Assumptions 1-4, the progress of task $s$’s global parameters at round $t+1$, $\overline{\textbf{w}}_{s}^{(t+1)}$, towards the optimal parameters $\textbf{w}^{*}$: $\mathbb{E}\left[\|\overline{\textbf{w}}^{(t+1)}_{s}-\textbf{w}_{s}^{*}\|^{2}\right]$ is upper-bounded by
$\left[1-\eta_{t}\mu(1+\frac{3\overline{\rho}^{\alpha}}{8})\right]\mathbb{E}[\|\overline{\textbf{w}}^{(t)}-\textbf{w}^{*}_{s}\|^{2}]+2\eta_{t}\Gamma_{s}(\tilde{\rho}^{\alpha}-\overline{\rho}^{\alpha})$
$+\eta_{t}^{2}\sigma^{2}\sum_{Sel(t)\in\mathbb{S}(t)}\sum_{k\in Sel(t)}B_{Sel(t)}^{s}(\alpha)\left(\frac{p_{k}}{\sum_{k^{\prime}\in Sel(t)}p_{k’}}\right)^{2}$
@@
Proof Sketch: We adapted the outline of [22]’s proof to incorporate our distinct client selection scheme, which unlike [22]’s proof requires explicitly quantifying the impact of $Sel(t)$ and $B^{s}_{Sel}$ (Eq. 7) through various bounds on the loss function gradients. In the rest of our proof, we add and subtract multiple expressions involving the gradients, bound the various terms, and incorporate Lemma 3’s result.
In tracking the impact of our client-task allocation on the convergence in Theorem 4’s proof, we also reveal how the convergence bound in (4) depends on our client-task allocation algorithm, in particular the parameter $\alpha$ and the current task losses. This result not only helps us to prove the convergence of our FedFairMMFL training algorithm, but also provides additional justification for our client-task allocation strategy: Mathematically, tasks that are further from the optimum (i.e., with a larger $\mathbb{E}[\|\overline{\textbf{w}}^{(t)}-\textbf{w}^{*}_{s}\|^{2}$) should be allocated clients to reduce the right-hand side of (4) to equalize the progress across all tasks, which agrees with our algorithm. In Corollary 5, we show that under certain conditions on $p_{k}$, as $\alpha$ increases, the term with $B^{s}_{Sel}(\alpha)$ decreases, hence controlling the task’s progress towards its optimum.
@@
Suppose that all of the $p_{k}=\frac{1}{K}$ and consider task $s$ that has the highest loss $f_{s}\left(\textbf{w}_{s}(t+1)\right)$. Then as $\alpha$ increases, the term $\eta_{t}^{2}\sigma^{2}\sum_{Sel\in\mathbb{S}(t)}\sum_{k\in Sel}B_{Sel}^{s}(\alpha)\left(\frac{p_{k}}{\sum_{k^{\prime}\in S(t)}p_{k’}}\right)^{2}$ in the right-hand side of (4) decreases.

### v2

@@ IV-A Fairness Guarantees
We first analyze the optimal solution to our fairness objective (3) to show that it reduces the variation in training accuracy, which intuitively improves fairness across tasks. Firstly, we let $G_{\alpha}$ represent the optimal value of our MMFL fairness objective (3), given that each client only trains one task each global epoch.
$G_{\alpha}=\min_{{\color[rgb]{0,0,0}\mathbf{w}_{1},...\mathbf{w}_{S}}}\sum^{S}_{s=1}f^{\alpha}_{s}(\textbf{w}_{s})\quad\text{s.t.}\sum_{s}a_{k,s}(t)\leq 1\ \forall k\ \forall t,$ (7)
where $a_{k,s}(t)=1$ indicates that client $k$ trains task $s$ during epoch $t$, and $\sum_{s}a_{k,s}(t)\leq 1\ \forall k\ \forall t$ expresses the constraint that each client trains 1 task each epoch at most. We let $(\mathbf{w}^{*}_{1_{\alpha}},...\mathbf{w}^{*}_{S_{\alpha}})$ denote the global optimal solution for (7) when parameter $\alpha$ is used.
To motivate the choice of $\alpha$, we show that for $\alpha=1$ and $2$, a higher $\alpha$ puts more emphasis on fairness: we relate this $\alpha$-fairness metric to the well-known variance and cosine similarity metrics for similarity across task accuracies. Note that the specific case of $\alpha=2$ is equivalent to the square operation on $f_{s}$.
@@
$\alpha=2$ yields a smaller variance in the task losses than $\alpha=1$ does:
$\textbf{Var}(f_{1}(\textbf{w}_{1_{2}}^{*}),...,f_{S}(\textbf{w}_{S_{2}}^{*}))\leq\textbf{Var}(f_{1}(\textbf{w}_{1_{1}}^{*}),...,f_{S}(\textbf{w}_{S_{1}}^{*})).$
The left hand side of this equation refers to the variance in the task losses, at the optimal solution of $G_{\alpha}$, when $\alpha=2$. The right hand side of this equation refers to the variance in the task losses, at the optimal solution of $G_{\alpha}$, when $\alpha=1$.
Proof.
$\textbf{Var}(f_{1}(\mathbf{w}_{1_{2}}^{*}),...,f_{n}(\mathbf{w}_{n_{2}}^{*}))$ (8)
$=\frac{\sum_{s=1}^{S}f^{{\color[rgb]{0,0,0}2}}_{s}(\mathbf{w}_{s_{2}}^{*})}{S}-\bigg{(}\sum_{s=1}^{S}f^{{\color[rgb]{0,0,0}1}}_{s}(\mathbf{w}_{s_{2}}^{*})\bigg{)}^{2}$ (9)
$\leq\frac{\sum_{s=1}^{S}f^{{\color[rgb]{0,0,0}2}}_{s}(\mathbf{w}_{s_{1}}^{*})}{S}-\bigg{(}\sum_{s_{1}}^{S}f^{{\color[rgb]{0,0,0}1}}_{s}(\mathbf{w}_{s_{2}}^{*})\bigg{)}^{2}$ (10)
$\leq\frac{\sum_{s=1}^{S}f^{{\color[rgb]{0,0,0}2}}_{s}(\mathbf{w}_{s_{1}}^{*})}{S}-\bigg{(}\sum_{s=1}^{S}f^{{\color[rgb]{0,0,0}1}}_{s}(\mathbf{w}_{s_{1}}^{*})\bigg{)}^{2}.$ (11)
where the first inequality follows from the fact that $(\mathbf{w}^{*}_{1_{2}},...\mathbf{w}^{*}_{S_{2}})$ is the optimal solution of $G_{2}$ (Eq. (7)), which is $\min_{\mathbf{w}_{1},...\mathbf{w}_{S}}\bigg{\{}\bigg{(}\sum^{S}_{s=1}f^{{\color[rgb]{0,0,0}\alpha=2}}_{s}(\mathbf{w}_{s})\bigg{)},\text{s.t.}\sum_{s}a_{k,s}\leq 1\ \forall k\ \forall t\bigg{\}}$, and the second inequality follows from the fact that $(\mathbf{w}^{*}_{1_{1}},...\mathbf{w}^{*}_{S_{1}})$ is the optimal solution of $G_{1}$ (Eq. (7)), which is $\min_{\mathbf{w}_{1},...\mathbf{w}_{S}}\bigg{\{}\bigg{(}\sum^{S}_{s=1}f^{{\color[rgb]{0,0,0}\alpha=1}}_{s}(\mathbf{w}_{s})\bigg{)},\text{s.t.}\sum_{s}a_{k,s}\leq 1\ \forall k\ \forall t\bigg{\}}$. ∎
Next, we use the definition of cosine similarity to show that $\alpha=2$ results in a fairer solution than $\alpha=1$ (non-fairness). A higher cosine similarity between two quantities indicates that the quantities are more uniform in value.
@@
$\alpha=2$ results in a higher cosine similarity across tasks than $\alpha=1$, i.e.,
$\frac{\sum_{s=1}^{S}f_{s}(\textbf{w}_{s_{2}}^{*})}{\sqrt{\sum_{s=1}^{S}f_{s}^{{\color[rgb]{0,0,0}2}}(\textbf{w}_{s_{2}}^{*})}}\geq\frac{\sum_{s=1}^{S}f_{s}(\textbf{w}_{s_{1}}^{*})}{\sqrt{\sum_{s=1}^{S}f_{s}^{{\color[rgb]{0,0,0}2}}(\textbf{w}_{s_{1}}^{*})}}.$ (12)
Proof.
Since $(\mathbf{w}^{*}_{1_{1}},...\mathbf{w}^{*}_{S_{1}})$ is the optimal solution of $G_{1}=\min_{\mathbf{w}_{1},...\mathbf{w}_{S}}\bigg{\{}\bigg{(}\sum^{S}_{s=1}f^{{\color[rgb]{0,0,0}\alpha=1}}_{s}(\mathbf{w}_{s})\bigg{)},\text{s.t.}\sum_{s}a_{k,s}\leq 1\ \forall k\ \forall t\bigg{\}}$ according to (7), we have
$\sum_{s=1}^{S}f_{s}^{{\color[rgb]{0,0,0}1}}(\mathbf{w}_{s_{2}}^{*})\geq\sum_{s=1}^{S}f_{s}^{{\color[rgb]{0,0,0}1}}(\mathbf{w}_{s_{1}}^{*}).$ (13)
Since $(\mathbf{w}^{*}_{1_{2}},...\mathbf{w}^{*}_{S_{2}})$ is the optimal solution of $G_{2}=\min_{\mathbf{w}_{1},...\mathbf{w}_{S}}\bigg{\{}\bigg{(}\sum^{S}_{s=1}f^{{\color[rgb]{0,0,0}\alpha=2}}_{s}(\mathbf{w}_{s})\bigg{)},\text{s.t.}\sum_{s}a_{k,s}\leq 1\ \forall k\ \forall t\bigg{\}}$ according to (7), we have
$\sum_{t=1}^{S}f^{{\color[rgb]{0,0,0}2}}_{s}(\mathbf{w}_{s_{1}}^{*})\geq\sum_{t=1}^{S}f^{{\color[rgb]{0,0,0}2}}_{s}(\mathbf{w}_{s_{2}}^{*}).$ (14)
Therefore, we have
$\frac{\sum_{s=1}^{S}f_{s}(\textbf{w}_{s_{2}}^{*})}{\sqrt{\sum_{s=1}^{S}f_{s}^{2}(\textbf{w}_{s_{2}}^{*})}}\geq\frac{\sum_{s=1}^{S}f_{s}(\textbf{w}_{s_{1}}^{*})}{\sqrt{\sum_{s=1}^{S}f_{s}^{2}(\textbf{w}_{s_{1}}^{*})}}.$ (15)
@@ IV-B Convergence Guarantees
We first characterize $B_{Sel_{s}}^{s}(\alpha)$, the probability that a specific set of clients $Sel_{s}$ is selected for a specific task $s$ by our algorithm FedFairMMFL.
$B_{Sel_{s}}^{s}(\alpha)=\textstyle\left(\frac{f_{s}^{\alpha}(t)}{\sum_{s^{\prime}\in S}f_{s^{\prime}}^{\alpha}(t)}\right)^{|Sel_{s}|}\left(1-\frac{f_{s}^{\alpha}(t)}{\sum_{s^{\prime}\in S}f_{s^{\prime}}^{\alpha}(t)}\right)^{K-|Sel_{s}|}$ (16)
@@
$B_{Sel_{s}}^{s}(\alpha)$ follows the client-task allocation probability vector (Eq. (5)), and depends on the losses of all tasks $s\in S$, and our $\alpha$-fair parameter $\alpha$. Next we introduce $\Gamma_{s}$, the gap between the loss evaluated at the global optimum $\mathbf{w}_{s}^{*}$, and the loss evaluated at client $k$’s optimal $\mathbf{w}^{*}_{k,s}$, for task $s$, summed over all clients:
$\Gamma_{s}$ $=\sum_{k=1}^{K}p_{k,s}(F_{k,s}(\mathbf{w}_{s}^{*})-F_{k,s}(\mathbf{w}^{*}_{k,s}))$
$=f_{s}^{*}-\sum_{k=1}^{K}p_{k,s}F_{k,s}^{*}.$ (17)
We further characterise the selection skew $\rho^{\alpha}(B^{s}_{Sel_{s}},\mathbf{w}^{\prime})$. The selection skew is a ratio of the expected ‘local-global objective gap’ of the selected clients under FedFairMMFL over the ‘local-global objective gap’ of all $K$ clients, where $\textbf{w}^{\prime}$ is the point at which $F_{k,s}$ and $f_{s}$ are evaluated. This skew depends on the expectation of client allocation combinations, and hence on $B_{Sel_{s}}^{s}$ and the $\alpha$-fair parameter $\alpha$:
$\rho^{\alpha}(B^{s}_{Sel_{s}},\mathbf{w}^{\prime})$ (18)
$=\frac{\sum_{Sel_{s}\in\mathbb{S}}{\color[rgb]{0,0,0}B_{Sel_{s}}^{s}(\alpha)}\sum_{k\in Sel_{s}}{\color[rgb]{0,0,0}\frac{p_{k,s}(F_{k,s}(\textbf{w}^{\prime}_{s})-F_{k,s}^{*})}{\sum_{k^{\prime}\in Sel}p_{k’,s}}}}{f_{s}(\mathbf{w}^{\prime})-\sum_{k=1}^{K}p_{k,s}F_{k,s}^{*}}$
@@
We use $\overline{\rho}^{\alpha}$ and $\tilde{\rho}^{\alpha}$ respectively to denote lower and upper bounds on the selection skew throughout the training. Recall that $\tau$ denotes the number of SGD epochs in local training. We then introduce the following assumptions used in our convergence analysis, which are standard in the FL literature [28].
Assumption 1: $F_{1,s},..,F_{k,s}$ are L-smooth, i.e.
$F_{k,s}(\mathbf{v})\leq F_{k,s}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{k,s}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2},$ (19)
for all $\mathbf{v}$ and $\mathbf{w}$.
Assumption 2: $F_{1,s},..F_{k,s}$ are $\mu$-strongly convex, i.e.
$F_{k,s}(\mathbf{v})\geq F_{k,s}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{k,s}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2},$ (20)
for all $\mathbf{v}$ and $\mathbf{w}$. Note that $\mu\leq L$ from Assumption 1.
@@
Assumption 3: For all clients $k$, for the minibatch $\xi_{k,s}$ uniformly sampled at random, the stochastic gradient is unbiased:
$\mathbb{E}[g_{k,s}(\mathbf{w}_{k,s},\xi_{k,s})]=\nabla F_{k,s}(\textbf{w}_{k,s}),$ (21)
and its variance is bounded:
$\mathbb{E}\|g_{k,s}(\mathbf{w}_{k,s},\xi_{k,s})-\nabla F_{k,s}(\textbf{w}_{k,s})\|^{2}\leq\sigma^{2}.$ (22)
@@
Assumption 4: Bounded stochastic gradient. For all clients $k$, the expected square norm of the stochastic gradient is bounded:
$\mathbb{E}\|g_{k,s}(\textbf{w}_{k,s},\xi_{k,s})\|^{2}\leq G^{2}.$ (23)
Next, we introduce Lemma 3, which bounds the expected average discrepancy between the global model for task $s$, and client $k$’s model for task $s$, under client selection set $Sel_{s}$.
@@
The expected average discrepancy between the global model for task $s$, $\overline{\textbf{w}}^{(t)}_{s}$ and client $k$’s model for $s$, $\textbf{w}_{k,s}^{(t)}$, under selection $k\in{\color[rgb]{0,0,0}Sel_{s}(t)}$, can be bounded as
$\mathbb{E}[\sum_{k\in{\color[rgb]{0,0,0}Sel_{s}(t)}}p_{k,{\color[rgb]{0,0,0}Sel_{s}}}\|\overline{\textbf{w}}^{(t)}_{s}-\textbf{w}_{k,s}^{(t)}\|^{2}]$ (24)
$\leq 16\eta_{t}^{2}\tau^{2}G^{2}\sum_{{\color[rgb]{0,0,0}Sel_{s}(t)}\in\mathbb{S}(t)}B_{Sel_{s}(t)}^{s}(\alpha)P^{k,2}_{{\color[rgb]{0,0,0}Sel_{s}(t)}}$ (25)
where $P^{k,2}_{Sel_{s}(t)}=\sum_{\begin{subarray}{c}k\neq k^{\prime},\\ k,k^{\prime}\in{\color[rgb]{0,0,0}Sel_{s}(t)}\end{subarray}}p_{k,s}\frac{p_{k^{\prime},s}}{(\sum_{k^{{}^{\prime\prime}}\in{\color[rgb]{0,0,0}Sel_{s}(t)}}p_{k^{\prime\prime},s})^{2}}$.
To prove Lemma 3, we modified the proof in [28], to incorporate the probability that a specific set of clients $Sel$ is allocated to the task, under our client selection scheme, and we sum over all possible sets of clients. The full proof is available in our Technical Report [41].
We use the above result in Lemma 3, to help us prove the next theorem. In the next theorem, without loss of generality we show the convergence rate of a task in the MMFL setting under FedFairMMFL, and show its dependence on the relative loss levels of all tasks. When the prevailing loss level of task $s$ is higher, its convergence will be sped up under our algorithm.
@@
Suppose we use under a decaying learning rate $\eta_{t}=\frac{1}{\mu(t+\gamma)}$ at each round $t$. Under Assumptions 1-4, the progress of task $s$’s global parameters at round $t+1$, $\overline{\textbf{w}}_{s}^{(t+1)}$, towards the optimal parameters $\textbf{w}^{*}$: $\mathbb{E}\left[\|\overline{\textbf{w}}^{(t+1)}_{s}-\textbf{w}_{s}^{*}\|^{2}\right]$ is upper-bounded by
$\left[1-\eta_{t}\mu(1+\frac{3\overline{\rho}^{\alpha}}{8})\right]\mathbb{E}[\|\overline{\textbf{w}}^{(t)}_{s}-\textbf{w}^{*}_{s}\|^{2}]+2\eta_{t}\Gamma_{s}(\tilde{\rho}^{\alpha}-\overline{\rho}^{\alpha})+$
$\eta_{t}^{2}\sigma^{2}\sum_{{\color[rgb]{0,0,0}Sel_{s}(t)}\in\mathbb{S}(t)}\sum_{k\in{\color[rgb]{0,0,0}Sel_{s}(t)}}B_{{\color[rgb]{0,0,0}Sel_{s}(t)}}^{s}(\alpha)\left(\frac{p_{k,s}}{\sum_{k^{\prime}\in{\color[rgb]{0,0,0}Sel_{s}(t)}}p_{k’,s}}\right)^{2}$
@@
Proof Sketch: We modified [28]’s proof to incorporate our distinct client selection scheme, which unlike [28]’s proof requires explicitly quantifying the impact of ${\color[rgb]{0,0,0}Sel_{s}(t)}$ and $B^{s}_{{\color[rgb]{0,0,0}Sel_{s}}}$ (Eq. 16) through various bounds on the loss function gradients. In the rest of our proof, we add and subtract multiple expressions involving the gradients, bound the various terms, and incorporate Lemma 3’s result. The full proof is available in our Technical Report [41].
In tracking the impact of our client-task allocation on the convergence in Theorem 4’s proof, we also reveal how the convergence bound in (4) depends on our client-task allocation algorithm, in particular the parameter $\alpha$ and the current task losses. This result not only helps us to prove the convergence of our FedFairMMFL training algorithm, but also provides additional justification for our client-task allocation strategy: Mathematically, tasks that are further from the optimum (i.e., with a larger $\mathbb{E}[\|\overline{\textbf{w}}^{(t)}-\textbf{w}^{*}_{s}\|^{2}$) should be allocated clients to reduce the right-hand side of (4) to equalize the progress across all tasks, which agrees with our algorithm. In Corollary 5, we show that under certain conditions on $p_{k,s}$, as $\alpha$ increases, the term with $B^{s}_{{\color[rgb]{0,0,0}Sel_{s}}}(\alpha)$ decreases, hence controlling the task’s progress towards its optimum.
@@
Suppose that all of the $p_{k,s}=\frac{1}{K}$ and consider task $s$ that has the highest loss $f_{s}\left(\textbf{w}_{s}(t+1)\right)$. Then as $\alpha$ increases, the term
${\color[rgb]{0,0,0}\eta_{t}^{2}\sigma^{2}\sum_{{\color[rgb]{0,0,0}Sel_{s}}\in\mathbb{S}(t)}\sum_{k\in{\color[rgb]{0,0,0}Sel_{s}(t)}}B_{{\color[rgb]{0,0,0}Sel_{s}}}^{s}(\alpha)\left(\frac{p_{k,s}}{\sum_{k^{\prime}\in{\color[rgb]{0,0,0}Sel_{s}(t)}}p_{k’,s}}\right)^{2}}$ (27)
in the right-hand side of (4) decreases.